The length of a rectangle is five more than twice the width. If the perimeter is 70 feet, find the length?
Answer by texttutoring(324)

Let l=length and w=width
The formula for perimeter is P=2l+2w and we know that l=2w+5

70=2l+2w

Now replace the l with 2w+5

70=2(2w+5) + 2w
70=4w+10+2w
70=6w+10
60=6w
w=10

l=2w+5
l=2(10)+5
l=25

So the width is 10, and the length is 25.